# Poincaré inequality and Quantitative De Giorgi method for hypoelliptic operators

### v2

@@ Abstract.
We propose a systematic elementary approach based on trajectories to prove Poincaré inequalities for hypoelliptic equations with an arbitrary number of Hörmander commutators, both in the local and in the non-local case. Our method generalises and simplifies the method introduced in [8] in the local case with one commutator, and extended in [13] to the non-local case with one commutator. It draws inspiration from the paper [17], although we use different trajectories. We deduce Harnack inequalities and Hölder regularity along the line of the De Giorgi method. Our results recover those in [2] in the local case, and are new in the non-local case.
AL is supported by the Cambridge Trust. CM is supported by ERC grant MAFRAN
@@ 1.1. The question at hand
This paper is devoted to the regularity theory of hypoelliptic equations with rough coefficients. We consider equations with an arbitrary number of commutators to obtain regularity in all directions, in the sense of Hörmander [9], and whose diffusive part is either local (second-order) or non-local (fractional order). Beyond their structural intrinsic importance, such equations naturally appear in financial models (see Section 1.5).
We consider an unknown function $f=f(x,t)$ satisfying equations of the form
@@
where ${\mathcal{B}}$ is a first-order differential operator and ${\mathcal{A}}$ is a differential, or fractional differential operator in some of the directions $x^{(0)}$ of $x$. We assume no regularity on the coefficients of the diffusive part ${\mathcal{A}}$. Then provided that ${\mathcal{A}}$ and ${\mathcal{B}}$ satisfy the Hörmander commutator condition [9], one expects the following Poincaré inequality to hold for weak non-negative sub-solutions:
$\int_{Q_{+}}\left(f-\frac{1}{|Q_{-}|}\int_{Q_{-}}f\right)_{+}\lesssim\int_{\Omega}|{\mathcal{A}}f|$
where $Q^{-},Q^{+}$ are two disjoint cylinders in the open domain $\Omega$ so that $Q^{+}$ is a time-translation of $Q^{-}$ into the future. In this paper we prove such an inequality with a quantitative control on the constant for a large class of equations of the type (1.1). The main novelty is the method itself, but the result is also new when $\mathcal{A}$ is non-local and more than one commutator is involved.
The proof is based on the combination of a method introduced in [8] and the explicit construction of appropriate new trajectories inspired partly from [17]. The Poincaré inequality is a powerful integral way to measure the oscillation of the sub-solution, and its use in this context goes back to [16, 15]. We deduce quantitatively the De Giorgi lemmas, the Hölder regularity for weak solutions, and the (weak) Harnack inequality for sub-solutions following the methodology recently developed in [8, 13].
@@ 1.2. Examples of equations
When the number of commutators $\kappa\geq 2$, the prototypical local equation is
@@
We denote $v=x^{(0)}$ for notational consistency.
In the fractional (non-local) case and with $\kappa=1$ commutators, the prototypical equation is the following fractional Kolmogorov equation
@@
on the unknown $f$ which depends on $x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},x^{(0)}\in{\mathbb{R}}^{d}$, $t\in{\mathbb{R}}$, and where the function $a$ is measurable and so that, for some $\Lambda\geq 1$,
(1.9) $\forall\,(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)\in{\mathbb{R}}^{(\kappa+1)d+1},\quad\Lambda^{-1}\leq a(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)\leq\Lambda.$
In fact, our results still apply when the fractional diffusion operator on the right hand side of (1.6) and (1.8) is replaced by the more general nonlocal operator
$\mathcal{L}f\left(x^{(\kappa)},\dots,x^{(1)},x^{(0)},t\right)$
$=\int_{{\mathbb{R}}^{d}}\left[f\left(x^{(\kappa)},\dots,x^{(1)},w,t\right)-f\left(x^{(\kappa)},\dots,x^{(1)},x^{(0)},t\right)\right]K\left(x^{(\kappa)},\dots,x^{(1)},x^{(0)},w,t\right){\,\mathrm{d}}w$
@@ 1.3. General assumptions (H)
We denote the variable $z:=(x,t)\in{\mathbb{R}}^{N+1}$ with $t\in{\mathbb{R}}$ and $x=(x^{(\kappa)},\dots,x^{(1)},x^{(0)})$ with each $x^{(i)}\in{\mathbb{R}}^{d_{i}}$, $d_{i}\in{\mathbb{N}}^{*}$ and $i\in\{0,\dots,\kappa\}$. We assume that $d_{0}\geq d_{1}\geq\dots\geq d_{\kappa}\geq 1$ and denote $N:=d_{0}+\cdots+d_{\kappa}$.
@@
with $d_{i}\times d_{i-1}$ blocks ${\mathsf{B}}_{i}$ which are constant matrices of rank $d_{i}$.
The operator ${\mathcal{A}}$ is a linear operator acting only on the variable $x^{(0)}$ and $\beta$-homogeneous in this variable with $\beta\in(0,1]$ and it satisfies the following pointwise control from above and integral control from below, given $\varphi\in C^{\infty}_{c}({\mathbb{R}}^{N})$:
(1.13) $\|{\mathcal{A}}\varphi\|_{\infty}\lesssim\|\nabla_{(0)}\varphi\|_{L^{\infty}}^{\beta}\|\varphi\|_{L^{\infty}}^{1-\beta},\quad\int\left|{\mathcal{A}}\varphi\right|{\,\mathrm{d}}z\gtrsim\int\left|(-\Delta_{v})^{\beta/2}\varphi\right|{\,\mathrm{d}}z.$
@@
This transformation leaves the operator $\mathcal{T}$ invariant: $(\mathcal{T}f)(\tilde{z}\circ z)=\mathcal{T}[f(\tilde{z}\circ\cdot)](z)$.
The scaling properties of ${\mathcal{A}}$ and ${\mathcal{B}}$ naturally define the hypoelliptic dilations for $r>0$:
(1.15) $\delta_{r}:=\textrm{diag}\left(r^{1+2\kappa\beta}\mathbb{I}_{d_{\kappa}},\dots,r^{1+2\beta}\mathbb{I}_{d_{1}},r\mathbb{I}_{d_{0}},r^{2\beta}\right)$
@@
The (local) higher-order Kolmogorov equation (1.4) corresponds to $d_{i}=d$ for all $i$, $(x^{(\kappa)},\dots,x^{(1)},x^{(0)})\in{\mathbb{R}}^{(\kappa+1)d}$, ${\mathcal{A}}:=\sqrt{A}\nabla_{(0)}$ with a $d\times d$-matrix $A=A(x^{(\kappa)},\dots,x^{(1)},x^{(0)},t)$ measurable symmetric so that $A\in[\Lambda^{-1},\Lambda]$, and the matrix
@@ 1.4. Main result
A function $f$ is a super-solution of (1.1) if $-f$ is a sub-solution.
@@
Consider operators ${\mathcal{A}},{\mathcal{B}}$ satisfying the assumptions (H). Let $R>0$ be sufficiently large, depending on the number of commutators $\kappa$. Let $\Omega=B_{R}\times\dots\times B_{R}\times(-5,0)$ and suppose $f$ is a weak non-negative sub-solution to (1.1) in $\Omega$. Then
@@
where $Q^{+}:=Q_{1}\subset\Omega$ denotes the future cylinder, and $Q^{-}=B_{1}\times\ldots\times B_{1}\times(-5,-4]\subset\Omega$ denotes the past cylinder (see Figure 1). Moreover, if $\beta=1$, that is in the local case, $\tilde{\Omega}=\Omega$, whereas, if $\beta\in(0,1)$, that is in the non-local case, $\tilde{\Omega}={\mathbb{R}}^{d_{0}}\times B_{R}\times\dots\times B_{R}\times(-5,0)$. The constant is universal and depends on $\kappa$, $\beta$, $d_{0},\dots,d_{\kappa}$, ${\mathsf{B}}$, $R$ and $\Lambda$.
[…]
@@
    The radius $R$ of the ambient space $Q_{R}$ is large enough such that the trajectories constructed in the proof (see (2.2)) do not exit $Q_{R}$. Since all maps are affine, we know that $R<+\infty$. One could try to optimise the size of the ambient space by optimising the control function used in the construction of the trajectories.
@@
    It is easy to include a source term $S$ in the right hand side of (1.1). Then the $L^{1}$ norm of $S$ on $\Omega$ should appear on the right hand side of (1.18).
@@
    $\int_{Q^{+}}\big{(}f-\langle f\rangle_{Q^{-}}\big{)}_{+}^{p}{\,\mathrm{d}}z\lesssim\int_{\Omega}\big{|}{\mathcal{A}}f\big{|}^{p}{\,\mathrm{d}}z.$
    In fact, it is even possible to obtain directly a gain of integrability for $f$ inside the Poincaré by our proof, see the related comments in Remark 8.
@@
    (1.19) $\int_{\Omega}\left(f-\int_{\Omega}f\psi_{-}\right)\psi_{+}{\,\mathrm{d}}z\lesssim_{\psi_{\pm}}\int_{\Omega}\big{|}{\mathcal{A}}f\big{|}{\,\mathrm{d}}z$
@@
    In such a mollified formulation, the Poincaré inequality in the local case would follow from the following “wholespace hypoelliptic Bogovoskiǐ result” (in the spirit of [4]): given $g=\psi_{+}-\psi_{-}$ with $\psi_{\pm}$ as above, construct a vector field $\mathbf{F}:=(F_{T},F_{(0)})$ with $F_{T}$ one-dimensional and $F_{(0)}$ $d_{0}$-dimensional, so that
    ${\mathcal{T}}^{*}F_{T}-\nabla_{(0)}\cdot F_{(0)}=g\quad\text{ and }\quad F_{T}\geq 0\quad\text{ and }\quad\|\nabla_{(0)}F_{T}\|_{\infty}+\|F_{(0)}\|_{\infty}\lesssim_{g}1.$
@@
    It is also possible to prove the local case of our Poincaré inequality through a “parabolic wholespace Bogovoskiǐ inequality” in $(x^{(\kappa)},\dots,x^{(1)},t)$ (excluding the $x^{(0)}$ variable, i.e. the velocity in case of a Kolmogorov equation), by following the arguments in [6]. This parabolic wholespace Bogovoskiǐ inequality is an easy generalisation of the Bogovoskiǐ inequality, where the explicit representation of Bogovoskiǐ is used to show that the first coordinate of the vector field solving the divergence problem is controlled below, due to the form of $g$ given in the previous point (6).
Our second main result is the weak Harnack inequality
@@
Let the operators ${\mathcal{A}},{\mathcal{B}}$ be such that ${\mathcal{B}}$ is given by (H) and ${\mathcal{A}}$ is more specifically given by
@@
Let $f$ be a super-solution to (1.1) on $Q_{1}$ with $0\leq f\leq 1$ in $Q_{1}$ (local case) or on ${\mathbb{R}}^{d}\times B_{1}^{\otimes\kappa}\times[-1,0]$ with $0\leq f\leq 1$ (non-local case). Then there is $C>0$ and $\zeta>0$ depending on $\beta,N,\kappa,\Lambda$ such that for $r_{0}>0$ sufficiently small the Weak Harnack inequality is satisfied:
@@
where $\tilde{Q}_{\frac{r_{0}}{2}}^{-}:=Q_{\frac{r_{0}}{2}}\big{(}(0,\dots,0,-\frac{5}{2}r_{0}^{2\beta}+\frac{1}{2}\big{(}\frac{r_{0}}{2}\big{)}^{2\beta})\big{)}\in{\mathbb{R}}^{N+1}$ (the local case is $\beta=1$).
@@
    In the local case the strong Harnack follows immediately from the weak Harnack inequality above and the gain of integrability (first lemma of De Giorgi). By contrast, the strong Harnack inequality does not follow as immediately in the non-local case, since the function values outside the domain affect the solution inside the domain. It is however possible to adapt the recent work [12, 14], which proves the strong Harnack inequality in the non-local case with one commutator, to our setting.

### v3

@@ Abstract.
We propose a systematic approach based on trajectories to prove a Poincaré inequality for weak non-negative sub-solutions to hypoelliptic equations with an arbitrary number of Hörmander commutators, both in the local and in the non-local case. As a consequence, we deduce the weak Harnack inequality and Hölder regularity along the line of the De Giorgi method.
AL is supported by the Cambridge Trust.
@@ 1.1. The question at hand
This paper is devoted to the regularity theory of hypoelliptic equations with rough coefficients. We consider equations with an arbitrary number of commutators in the sense of Hörmander [10], with a diffusive part that is either local, in which case the equation is of second order, or non-local, in which case we deal with an equation of fractional order. In either case though, the diffusion is degenerate, as it acts only along the direction of one vector field. Our aim is to recover regularity in all directions, by transferring regularity from the degenerate diffusivity onto all vector fields.
Concretely, we consider an unknown function $f=f(x,t)$ solving an equation of the form
@@
where $\mathcal{B}$} is a first-order differential operator and $\mathcal{A}$} is a, possibly fractional, differential operator. Here $t\in{\mathbb{R}}$ and $x=(x^{(\kappa)},\dots,x^{(1)},x^{(0)})$ with each $x^{(i)}\in{\mathbb{R}}^{d_{i}}$, $d_{i}\in{\mathbb{N}}^{*}$ and $i\in\{0,\dots,\kappa\}$ such that $N:=d_{0}+\cdots+d_{\kappa}$. The differential operator $\mathcal{A}$} acts merely in one direction of $x$, namely $x^{(0)}$, which makes up the diffusive part of the equation. Even though we assume no regularity on the coefficients of $\mathcal{A}$}, we still expect the diffusivity that stems from $\mathcal{A}$} to regularise the solution $f$ of (1.1) at least in the variable $x^{(0)}$. To recover a regularisation in all remaining directions of $x$, we require a so-called Hörmander’s commutator condition: the constant coefficient vector fields, and their commutators (defined as [X,Y]=XYYX[X,Y]=XY-YX, where $X$ and $Y$ are any two smooth vector fields) appearing in (1.1) are supposed to span the whole space. Any equation with this property is called hypoelliptic. In our notation, $\kappa$ denotes the number of commutators.
The aim of the current paper is to derive a Poincaré inequality for non-negative weak sub-solutions to (1.1), which exploits Hörmander’s hypoellipticity to deal with the degeneracy in the diffusive part:
(1.2) $\int_{Q_{+}}\left(f-\frac{1}{|Q_{-}|}\int_{Q_{-}}f\right)_{+}\lesssim\int_{\Omega}|{\mathcal{A}}f|,$
where $Q^{-},Q^{+}$ are two disjoint cylinders in the open domain $\Omega\subseteq\mathbb{R}^{N+1}$ so that $Q^{+}$ is a time-translation of $Q^{-}$ into the future. The symbol $\lesssim$ denotes that the inequality holds up to a universal constant $C$. The striking feature of (1.2) is that on the right hand side there appears merely a differential in the diffusive direction: it is crucial that no differential in other directions appears in order to deduce the weak Harnack inequality or Hölder regularity of solutions to (1.1).
In this paper we prove a Poincaré inequality with a quantitative control on the constant hidden in the symbol $\lesssim$ in (1.2) for a large class of equations of type (1.1). The Poincaré inequality is a powerful integral way to measure the oscillation of the sub-solution, and its use in this context goes back to [17, 16]. We use it to quantitatively to deduce the De Giorgi lemmas (and thus the Hölder regularity for weak solutions), and the (weak) Harnack inequality for super-solutions following the methodology recently developed in [8, 14]. The main novelty is the method itself, but the result is also new when $\mathcal{A}$ is a fractional differential operator in the case that more than one commutator is involved (i.e. $\kappa>1$).
The proof is based on a systematic construction of trajectories that encode the flow of the vector fields underlying (1.1), in such a way that they relate the future cylinder $Q^{+}$ to any point in the past cylinder $Q^{-}$. The use of trajectories is inspired from [8], where they consider the local case (second order equation) with one commutator $\kappa=1$. Their trajectories were combining the vector fields of the equation in a piecewise affine way; yet this was sub-optimal and they had an error term on the right hand side of (1.2). This approach was improved by L. Niebel and R. Zacher in [18], where they construct an ad-hoc smooth curve instead of a piecewise affine curve; however this was not sufficient to get rid of the error term that also appeared in [8]. Here we introduce an approach which picks up on ideas of both these papers: we construct a smooth trajectory as the solution of a control problem associated to the equation (1.1). We end up with a clean Poincaré inequality in the form (1.2) without any error terms on the right hand side.
@@ 1.2. Examples of equations
When $\kappa\geq 2$, the underlying geometry becomes more involuted, and the prototypical local equation is
@@
From now on, we denote $v=x^{(0)}$ for notational consistency.
In the fractional (non-local) case with $\kappa=1$ commutators, the prototypical equation is the following fractional Kolmogorov equation
@@
One should think about a divergence form equation of fractional order $2\beta\in(0,2)$ with rough diffusion coefficients. The fractional diffusion in velocity is the generator of a Lévy process. Under more general assumptions on the roughness of the coefficient $a$, this equation is a reformulation of the Boltzmann equation (without cutoff) in a specific conditional regime, up to a lower order source term. We will not specify this direction further in this work, since our focus is the abstraction and the development of a method that is robust enough to treat a wide class of operators that we believe can be generalised to specific models as needed. The robustness of the method, however, is best demonstrated within a class of equations that can be described under a notation that we hope to be readable.
@@
on the unknown $f$ which depends on $x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v\in{\mathbb{R}}^{d}$, $t\in{\mathbb{R}}$, and where the function $a$ is measurable and so that, for some $\Lambda\geq 1$,
It is more common for non-local equations to phrase the operator $\left(-\Delta_{v}\right)^{\frac{\beta}{2}}\cdot\left[a\left(-\Delta_{v}\right)^{\frac{\beta}{2}}f\right]$ in (1.7) as
$\mathcal{L}f\left(x^{(\kappa)},\dots,x^{(1)},v,t\right)$
$=\int_{{\mathbb{R}}^{d}}\left[f\left(x^{(\kappa)},\dots,x^{(1)},w,t\right)-f\left(x^{(\kappa)},\dots,x^{(1)},v,t\right)\right]K\left(x^{(\kappa)},\dots,x^{(1)},v,w,t\right){\,\mathrm{d}}w,$
@@ 1.3. General assumptions (H)
We denote the variable $z:=(x,t)\in{\mathbb{R}}^{N+1}$ with $t\in{\mathbb{R}}$ and $x=(x^{(\kappa)},\dots,x^{(1)},v)$ with each $x^{(i)}\in{\mathbb{R}}^{d_{i}}$, $d_{i}\in{\mathbb{N}}^{*}$ and $i\in\{1,\dots,\kappa\}$, and $v\in{\mathbb{R}}^{d_{0}}$. We assume that $d_{0}\geq d_{1}\geq\dots\geq d_{\kappa}\geq 1$ and denote $N:=d_{0}+\cdots+d_{\kappa}$.
@@
with $d_{i}\times d_{i-1}$ blocks ${\mathsf{B}}_{i}$ which are constant matrices of rank $d_{i}$ with $|{\mathsf{B}}|\leq\Lambda$.
The operator $\mathcal{A}$} is a linear operator acting only on the variable $v$ and $\beta$-homogeneous in this variable with $\beta\in(0,1]$ and it satisfies the following pointwise control from above and integral control from below, given $\varphi\in C^{\infty}_{c}({\mathbb{R}}^{N})$ such that $\|\varphi\|_{L^{\infty}}\geq 1$:
(1.14) $\|{\mathcal{A}}\varphi\|_{\infty}\lesssim\|\nabla_{v}\varphi\|_{L^{\infty}}^{\beta+\epsilon}\|\varphi\|_{L^{\infty}}^{1-\beta-\epsilon},\qquad\int\left|{\mathcal{A}}\varphi\right|{\,\mathrm{d}}z\gtrsim\int\left|(-\Delta_{v})^{\beta/2}\varphi\right|{\,\mathrm{d}}z,$
where $\epsilon=0$ for $\beta=1$, and $\epsilon\in(0,1-\beta)$ for $\beta\in(0,1)$. We point out that in case of $\beta=1$, we identify with a slight abuse of notation $(-\Delta_{v})^{\frac{1}{2}}$ with $\nabla_{v}$, such that we are not violating the incorrect inequality $\int\left|\nabla\varphi\right|{\,\mathrm{d}}z\gtrsim\int\left|(-\Delta_{v})^{1/2}\varphi\right|{\,\mathrm{d}}z$. In fact, in the local case where $\beta=1$, there is no need to introduce fractional derivatives at all, but in order to treat $\beta\in(0,1]$ we do so to ease the notation.
@@
This transformation leaves the operator $\mathcal{T}$ invariant: $(\mathcal{T}f)(\tilde{z}\circ z)=\mathcal{T}[f(\tilde{z}\circ\cdot)](z)$. When considering the principal part operator associated to (1.1), which is obtained by replacing $A$ in (1.5) with the identity matrix of dimension $d$ in the local case or by choosing $a$ in (1.9) as the identity in the non-local case, the related equation naturally defines scaling properties encoded by the following hypoelliptic dilations, defined for $r>0$
(1.16) $\delta_{r}:=\textrm{diag}\left(r^{1+2\kappa\beta}\mathbb{I}_{d_{\kappa}},\dots,r^{1+2\beta}\mathbb{I}_{d_{1}},r\mathbb{I}_{d_{0}},r^{2\beta}\right),$
@@
The (local) higher-order Kolmogorov equation (1.5) corresponds to $d_{i}=d$ for all $i$, $(x^{(\kappa)},\dots,x^{(1)},v)\in{\mathbb{R}}^{(\kappa+1)d}$, ${\mathcal{A}}:=\sqrt{A}\nabla_{v}$ with a $d\times d$-matrix $A=A(x^{(\kappa)},\dots,x^{(1)},v,t)$ measurable symmetric so that $A\in[\Lambda^{-1},\Lambda]$, and the matrix
@@ 1.4. Main result
A function $f$ is a super-solution of (1.1) if f-f is a sub-solution.
@@
Consider operators $\mathcal{A}$$\mathcal{B}$} satisfying the assumptions (H). Let $R>0$ be sufficiently large, depending on the number of commutators $\kappa$ and $\Lambda$. Let $Q_{R}:=\delta_{R}\left(Q_{1}\right)$ and suppose $f$ is a weak non-negative sub-solution to (1.1) in $\tilde{\Omega}$ where $\tilde{\Omega}=Q_{R}$ if $\beta=1$, that is in the local case, and $\tilde{\Omega}=B_{R}\times\dots\times B_{R}^{1+2(\kappa-1)\beta}\times{\mathbb{R}}^{d_{0}}\times(-5,0)$, if $\beta\in(0,1)$, that is in the non-local case. Then
@@
where $Q^{+}:=Q_{1}\subset\tilde{\Omega}$ denotes the future cylinder, and $Q^{-}=B_{1}\times\ldots\times B_{1}\times(-5,-4]\subset\tilde{\Omega}$ denotes the past cylinder (see Figure 1). The constant is universal and depends on $\kappa$, $\beta$, $d_{0},\dots,d_{\kappa}$, $\mathsf{B}$}, $R$ and $\Lambda$.
$\tilde{\Omega}$$^{+}QQ^{-}$
@@
    The radius $R$ of the ambient space $\tilde{\Omega}$ is large enough such that the trajectories constructed in the proof (see (2.2)) do not exit $\tilde{\Omega}$. Since all maps involved in the construction of such trajectories are affine, we know that $R<+\infty$. One could try to optimise the size of the ambient space by optimising the control function used in the construction of the trajectories.
@@
    It is easy to include a source term $S$ in the right hand side of (1.1). Then the $L^{1}$ norm of $S$ on $\tilde{\Omega}$ should appear on the right hand side of (1.19).
@@
    $\int_{Q^{+}}\big{(}f-\langle f\rangle_{Q^{-}}\big{)}_{+}^{p}{\,\mathrm{d}}z\lesssim\int_{\tilde{\Omega}}\big{|}{\mathcal{A}}f\big{|}^{p}{\,\mathrm{d}}z.$
    Furthermore, it is even possible to adapt our proof to obtain a gain of integrability for $f$, see the related comments in Remark 8.
@@
    (1.20) $\int_{\tilde{\Omega}}\left(f-\int_{\tilde{\Omega}}f\psi_{-}\right)\psi_{+}{\,\mathrm{d}}z\lesssim_{\psi_{\pm}}\int_{\tilde{\Omega}}\big{|}{\mathcal{A}}f\big{|}{\,\mathrm{d}}z$
@@
    In such a mollified formulation, the Poincaré inequality in the local case would follow from the following “wholespace hypoelliptic Bogovoskiǐ result” (in the spirit of [4]): given $g=\psi_{+}-\psi_{-}$ with $\psi_{\pm}$ as above, construct a vector field $\mathbf{F}:=(F_{T},F_{v})$ with $F_{T}$ one-dimensional and $F_{v}$ $d_{0}$-dimensional, so that
    ${\mathcal{T}}^{*}F_{T}-\nabla_{v}\cdot F_{v}=g\quad\text{ and }\quad F_{T}\geq 0\quad\text{ and }\quad\|\nabla_{v}F_{T}\|_{\infty}+\|F_{v}\|_{\infty}\lesssim_{g}1.$
@@
    It is also possible to prove the local case of our Poincaré inequality through a “parabolic wholespace Bogovoskiǐ inequality” in $(x^{(\kappa)},\dots,x^{(1)},t)$ (excluding the $x^{(0)}=v$ variable, i.e. the velocity in case of a Kolmogorov equation), by following the arguments in [6]. This parabolic wholespace Bogovoskiǐ inequality is an easy generalisation of the Bogovoskiǐ inequality, where the explicit representation of Bogovoskiǐ is used to show that the first coordinate of the vector field solving the divergence problem is controlled below, due to the form of $g$ given in the previous point (6).
It is possible to employ such a Poincaré inequality to prove a weak Harnack inequality for weak super-solutions to equation (1.1). We only sketch the proof of this here (see Section 4 for further details), since it follows from previous literature. Nevertheless, it is one of the main blocks to complete the De Giorgi-Nash-Moser theory for weak solutions, which consists in proving local quantitative regularity results (such as the a priori boundedness of weak solutions, the intermediate value lemma and the Harnack inequality) for weak solutions, and for this reason it is now stated as an independent result.
@@
Let the operators $\mathcal{A}$$\mathcal{B}$} be such that $\mathcal{B}$} is given by (H) and $\mathcal{A}$} is more specifically given by
@@
    ${\mathcal{A}}:=\sqrt{A}\nabla_{v}$ with $A\in[\Lambda^{-1},\Lambda]$ measurable symmetric matrix (local case),
@@
Let $f$ be a super-solution to (1.1) on $Q_{1}$, such that in the local case, $0\leq f$ in $Q_{1}$, and such that in the non-local case, $0\leq f\leq 1$ on $B_{1}^{\otimes\kappa}\ \times{\mathbb{R}}^{d_{0}}\times[-1,0]$. Then there is $C>0$ and $\zeta>0$ depending on $\beta,N,\kappa,\Lambda$ such that for $r_{0}>0$ sufficiently small the Weak Harnack inequality is satisfied:
@@
where $\tilde{Q}_{\frac{r_{0}}{2}}^{-}:=Q_{\frac{r_{0}}{2}}\big{(}0,\dots,0,-\frac{5}{2}r_{0}^{2\beta}+\frac{1}{2}\big{(}\frac{r_{0}}{2}\big{)}^{2\beta}\big{)}\in{\mathbb{R}}^{N+1}$ (the local case is $\beta=1$).
@@
    In the local case the strong Harnack follows immediately from the weak Harnack inequality above and the gain of integrability (first lemma of De Giorgi). By contrast, the strong Harnack inequality does not follow as immediately in the non-local case, since the function values outside the domain affect the solution inside the domain. It is however possible to adapt the recent work [1, 13, 15], which proves the strong Harnack inequality in the non-local case with $\kappa=1$, to our setting.
@@
  3. (3)
    The boundedness of $f$ in the non-local case almost everywhere in the $x^{(0)}=v$ direction is required to make sense of the equation (1.1).